Consider the quadratic function; f(x)=r^2 + (x+s)(x-s) here r,s are arbitary constants, find the min possible value for f and also the value of x for which the min is achieved.
my email address is jesssie87@yahoo.com.au Answer by stanbon(75887) 
 Consider the quadratic function; f(x)=r^2 + (x+s)(x-s) here r,s are arbitary constants, find the min possible value for f and also the value of x for which the min is achieved.
------------------------
f(x)=r^2 + (x+s)(x-s)
f(x) = x^2-s^2+r^2
a = 1 ; b = 0 ; c = (r^2-s^2)
----
The minimum value occurs when x = -b/(2a) = 0/(2) = 0
---
f(0) = r^2-s^2
----
Min: (0, r^2-s^2)
======================
Cheers,
Stan H.
================